### v9

@@ Abstract
Automatic breast cancer classification in histopathology images is crucial for precise diagnosis and treatment planning. Recently, classification approaches based on the ResNet architecture have gained popularity for significantly improving accuracy by using skip connections to mitigate vanishing gradient problems, thereby integrating low-level and high-level feature information. Nevertheless, the conventional ResNet architecture faces challenges such as data imbalance and limited interpretability, necessitating cross-domain knowledge and collaboration among medical experts. This study effectively addresses these challenges by introducing a novel method for breast cancer classification, the Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model. It integrates a pre-trained ResNet50 model with a lightweight attention mechanism, embedding an attention module in the fourth layer of ResNet50 and incorporating two fully connected layers with LeakyReLU and ReLU activation functions to enhance feature learning capabilities. The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. It was also compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50. The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets. Additionally, to address the interpretability challenges, we introduce the Dynamic Threshold Grad-CAM (DT Grad-CAM) method. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 not only ensures high classification performance but also offers enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis.
@@ I Introduction
Breast cancer has become a leading cause of mortality, particularly among women, with its incidence rising continuously, raising global health concerns [1, 2]. Early, accurate diagnosis and treatment are pivotal in decreasing the mortality rate associated with breast cancer. Yet, medical practices face challenges such as insufficient healthcare resources and human error in interpretation, which impede the timely and precise detection and diagnosis of breast cancer. Against this backdrop, intelligent diagnostic assistance technologies have emerged as innovative solutions in the battle against breast cancer. Recent advances in deep learning have revolutionized breast cancer diagnosis, surpassing the limitations of traditional classification and recognition algorithms by integrating feature extraction, selection, and classification into a unified step. Deep learning automates feature extraction and classification through neural networks, reducing redundancy and enhancing efficiency, thus significantly improving diagnostic accuracy and workflow [3, 4, 5, 6, 7]. A pivotal innovation in this domain is developing and applying Convolutional Neural Networks (CNNs), specialized deep learning models adept at processing data, such as medical images, with a grid-like topology. In 1998, LeCun et al. [8] introduced Convolutional Neural Networks (CNNs), which progressively extract abstract features from images through multiple levels of convolutional and pooling operations. This process enables CNNs to achieve classification based on these extracted features. CNNs comprise convolutional, pooling, and fully connected layers, collaboratively facilitating the robust classification of medical images. The convolutional layer identifies local image features; the pooling layer reduces data dimensions while retaining critical information; and the fully connected layer maps these features to output classes, as illustrated in Fig. 1. Despite deep learning’s significant achievements in medical image classification, conventional CNN architectures often struggle with imbalanced medical image data. This paper proposed a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model designed to overcome the challenge of imbalanced data in CNN architectures. This innovative approach efficiently manages uneven category distributions and enhances the detection of malignant breast cancer in medical images. The primary contributions of this work are summarized as follows:
(a) This study introduces a novel lightweight attention mechanism network that adaptively adjusts channel feature extraction to enhance feature representation. This approach optimizes the network’s ability to discern relevant features for breast cancer classification.
(b) Integrating the attention mechanism with ResNet50 [9] creates a feature-rich model, achieving more efficient feature extraction by combining the strengths of both techniques. Specifically, within the attention mechanism’s fully connected network component, LeakyReLU [10] and ReLU [11] activation functions are employed. LeakyReLU is chosen for its capability to alleviate the issue of vanishing gradients; ReLU is used to improve the representation of image features, facilitating the capture of multi-scale spatial patterns more effectively.
(c) The paper tests images at various magnifications, demonstrating the network’s scale invariance and superior performance over other contemporary methods.
(d) The model undergoes comprehensive evaluation, focusing on analyzing loss. The findings indicate stable performance during the later stages of training, effectively minimizing overfitting and demonstrating strong generalization capabilities for new data.
@@ III-A Transfer ResNet50
TABLE II: Comparison between DALAResNet50 model and standard ResNet model
Feature DALAResNet50 model Standard ResNet model
Adaptive Average Pooling Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding. Not available
Channel Recalibration A two-layer fully connected network for channel recalibration optimize ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification. Not available
Activation Functions The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, nonzero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers. This combination maintains the model’s learning capability, potentially enhancing the training process. ReLU only
Element-wise multiplication Element-wise multiplication, as denoted by ’Scale’ in Fig. 3, enhance ”Feature Adaptability” by allowing dynamic, nonlinear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information. Not available
@@ III-B3 DALAResNet50 model optimization
Fig. 3 illustrates the flowcharts of the standard ResNet50 approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. Table 2 details a comprehensive comparative analysis of these differences. Fig. 4 depicts the visualization of the DALAResNet50 approach.
@@ IV-A BreakHis Dataset
Since the BreakHis dataset is a publicly available dataset, this study does not require ethical approval. Additionally, publicly available datasets are typically subject to stringent quality control measures; therefore, this study does not separately consider image quality issues.
40x
100x
200x
400x
Fig. 5: Images of breast cancer pathology at four magniations
TABLE III: BreakHis dataset: benign and malignant image distribution across magnification factors
@@
TABLE IV: Dataset split for experimental task
Split Percentage Purpose
Training 60% Model Training
Validation 10% Hyperparameter Tuning
Testing 30% Model Evaluation
TABLE V: Experimental configuration
Parameter Value
Batch Size 32
Epochs 50
Learning Rate 0.0001
Dropout 0.25
@@ IV-C Experiment implementation
TABLE VI: Classification results of 9 network models
@@
TABLE VII: The count of parameters in the network model and the time it takes to converge.
@@ IV-E Experiment metrics
Table 6 showcases the model’s performance on the test set at various image resolutions (40X, 100X, 200X, and 400X). Across these resolutions, the model consistently outperforms nine other popular networks regarding accuracy, F1 score, IBA, and GMean. This consistency underlines the network’s significant advantage and underscores its potential for medical image classification tasks.
Table 7 compares the model against eight others regarding parameter count and convergence speed. While DenseNet-121 has the fewest parameters, the proposed model achieves the shortest convergence times, exhibiting the fastest average speed across all resolutions (40X, 100X, 200X, and 400X). This further reinforces the proposed model’s efficiency and strong potential for medical image classification, highlighting superior performance and effectiveness.
@@ IV-F Ablation study
In this ablation study, we compared the performance of the DALAResNet50 model and the standard ResNet50 model for benign and malignant classification at 40x, 100x, 200x, and 400x magnifications. As shown in Table 6, the DALAResNet50 model outperformed the standard ResNet50 model across all metrics, including accuracy, precision, recall, F1 score, IBA, and GMean. Additionally, we provided confusion matrices for both models on the test set, illustrated in Fig. 6. The results demonstrated that the DALAResNet50 model exhibited higher sensitivity and specificity at all magnifications, significantly reducing the number of false positives and false negatives. These findings suggest that the DALAResNet50 model surpasses the standard ResNet50 model in terms of classification accuracy and reliability.
@@ 1. Grad-CAM Calculation
2. Otsu Threshold Calculation
where:
  • $\sigma_{B}^{2}(T)$ is the between-class variance for threshold $T$.
@@ V-B Comparison between DT Grad-CAM and Grad-CAM Visualization
TABLE VIII: Comparison between DT Grad-CAM and Grad-CAM
@@ VII DALAResNet50’s Future Therapeutic Impact and Integration into Existing Diagnostic Processes
Overall, the DALAResNet50 model, combined with DT Grad-CAM’s interpretability, holds great promise for transforming breast cancer diagnosis and treatment. Its integration into existing diagnostic workflows can lead to more precise, reliable, and efficient medical care, ultimately enhancing patient outcomes and advancing the field of histopathology.
@@ VIII Summary and conclusions
This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets. The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper’s model performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. Furthermore, this study introduces the DT Grad-CAM method to address the interpretability challenges often associated with deep learning models. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, which provides clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 ensures high classification performance and enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis. In future research, more investigations into breast multiclassification problems are planned.

### v10

@@ Abstract
Automatic breast cancer classification in histopathology images is crucial for precise diagnosis and treatment planning. Recently, classification approaches based on the ResNet architecture have gained popularity for significantly improving accuracy by using skip connections to mitigate vanishing gradient problems, thereby integrating low-level and high-level feature information. Nevertheless, the conventional ResNet architecture faces challenges such as data imbalance and limited interpretability, necessitating cross-domain knowledge and collaboration among medical experts. This study effectively addresses these challenges by introducing a novel method for breast cancer classification, the Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model. It integrates a pre-trained ResNet50 model with a lightweight attention mechanism, embedding an attention module in the fourth layer of ResNet50 and incorporating two fully connected layers with LeakyReLU and ReLU activation functions to enhance feature learning capabilities. The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. It was also tested on the four-class classification BACH dataset and the three-class classification Mini-DDSM dataset, where it also performed excellently. Additionally, the model was compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50. The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets, proving the model’s wide applicability. To address the interpretability challenges, we introduce the Dynamic Threshold Grad-CAM (DT Grad-CAM) method. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 not only ensures high classification performance but also offers enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis.
@@ I Introduction
Breast cancer has become a leading cause of mortality, particularly among women, with its incidence rising continuously, raising global health concerns [1, 2]. Early, accurate diagnosis and treatment are pivotal in decreasing the mortality rate associated with breast cancer. Yet, medical practices face challenges such as insufficient healthcare resources and human error in interpretation, which impede the timely and precise detection and diagnosis of breast cancer. Against this backdrop, intelligent diagnostic assistance technologies have emerged as innovative solutions in the battle against breast cancer. Recent advances in deep learning have revolutionized breast cancer diagnosis, surpassing the limitations of traditional classification and recognition algorithms by integrating feature extraction, selection, and classification into a unified step. Deep learning automates feature extraction and classification through neural networks, reducing redundancy and enhancing efficiency, thus significantly improving diagnostic accuracy and workflow [3, 4, 5, 6, 7]. A pivotal innovation in this domain is developing and applying Convolutional Neural Networks (CNNs), specialized deep learning models adept at processing data, such as medical images, with a grid-like topology. In 1998, LeCun et al. [8] introduced Convolutional Neural Networks (CNNs), which progressively extract abstract features from images through multiple levels of convolutional and pooling operations. This process enables CNNs to achieve classification based on these extracted features. CNNs comprise convolutional, pooling, and fully connected layers, collaboratively facilitating the robust classification of medical images. The convolutional layer identifies local image features; the pooling layer reduces data dimensions while retaining critical information; and the fully connected layer maps these features to output classes. Despite deep learning’s significant achievements in medical image classification, conventional CNN architectures often struggle with imbalanced medical image data. This paper proposed a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model designed to overcome the challenge of imbalanced data in CNN architectures. This innovative approach efficiently manages uneven category distributions and enhances the detection of malignant breast cancer in medical images. The primary contributions of this work are summarized as follows:
(a) This study introduces a novel lightweight attention mechanism network that adaptively adjusts channel feature extraction to enhance feature representation. This method optimizes the network’s ability to recognize features relevant to breast cancer classification. Performance was evaluated on the BreakHis dataset for binary classification, the BACH dataset for four-class classification, and the Mini-DDSM dataset for three-class classification. Evaluation metrics included accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of the model’s performance compared to existing techniques. Experimental results highlighted significant advantages, especially in addressing the challenges of imbalanced breast cancer datasets. The model demonstrated improved robustness, generalization, faster convergence during training, and effective prevention of overfitting.
(b) This paper introduces a DT Grad-CAM method to address the interpretability challenges commonly associated with deep learning models. DT Grad-CAM enhances the traditional Grad-CAM by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of important regions. This approach improves the clarity and relative prominence of highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the key features contributing to the classification.
@@ III-A Transfer ResNet50
40x
100x
200x
400x
Fig. 4: Images of breast cancer pathology at four magniations
@@ III-B3 DALAResNet50 model optimization
Fig. 2 illustrates the flowcharts of the standard ResNet50 approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. Fig. 3 depicts the visualization of the DALAResNet50 approach.
@@ IV-A1 BreakHis Dataset
Since the BreakHis dataset is publicly available, this study does not require ethical approval. Additionally, publicly available datasets are typically subject to stringent quality control measures; therefore, this study does not consider image quality issues separately.
IV-A2 BACH Dataset
The BACH (Breast Cancer Histology Images) dataset [63] is a significant breast cancer histopathology image dataset provided by the ICIAR2018 Grand Challenge. This dataset includes 400 HE-stained histopathology images of breast cancer tissue, divided into four categories: Normal, Benign, In-situ carcinoma, and Invasive carcinoma, with 100 images in each category. All images have a resolution of 2048 × 1536 pixels, As shown in Table 3. The BACH dataset is widely used for training and testing machine learning and deep learning models to develop and evaluate breast cancer detection and classification algorithms. It provides high-quality annotation information, making it an essential resource in breast cancer research.
IV-A3 Mini-DDSM Dataset
The Mini-DDSM (Mini Digital Database for Screening Mammography) [62] is a reduced version of the DDSM dataset, primarily used for research and development in breast cancer screening. This dataset contains approximately 2000 mammography images, as shown in Table 4. a subset of the DDSM database that retains critical pathological features. The photos in the Mini-DDSM dataset are annotated with detailed information about lesions, including the location of masses and calcifications and their pathological diagnoses.
TABLE II: BreakHis dataset: benign and malignant image distribution across magnification factors
@@ IV-A3 Mini-DDSM Dataset
TABLE III: BACH Dataset Details
Category Number of Images Image Size
Normal 100 2048 × 1536
Benign 100 2048 × 1536
In-situ carcinoma 100 2048 × 1536
Invasive carcinoma 100 2048 × 1536
Total 400 2048 × 1536
TABLE IV: Detailed Information of the Mini-DDSM Dataset
Category Number of Images Image Size / Format
Normal Images 500 Variable
Benign Images 700 Variable
Malignant Images 800 Variable
Total Images 2000 Variable
@@ IV-C Experiment implementation
TABLE V: Performance of Network Models on BreakHis Dataset (Two-Class Classification)
@@
TABLE VI: Performance of Network Models on BACH Dataset (Four-Class Classification)
Network Models Accuracy F1 IBA GMean
SEResNet50 [39] 0.770 0.743/0.77 0.583 0.553
DensNet121 [40] 0.703, 0.680/0.703 0.458 0.438
VGG16 [41] 0.788 0.781 /0.788 0.652 0.637
VGG16Inception [42] 0.727 0.864/0.727 0.650 0.654
ViT [43] 0.746 0.734/ 0.745 0.580 0.555
Swin-Transformer [44] 0.748 0.556/0.745 0.500 0.436
Dinov2_Vitb14 [45] 0.856 0.730/0.855 0.500 0.353
ResNet50 [9] 0.855 0.838 /0.855 0.664 0.635
DALAResNet50 0.891 0.891/0.885 0.748 0.734
TABLE VII: Performance of Network Models on Mini-DDSM Dataset (Three-Class Classification)
Network Models Accuracy F1 IBA GMean
SEResNet50 [39] 0.7006 0.699/0.701 0.709 0.768
DensNet121 [40] 0.6301 , 0.627/0.63 0.642 0.716
VGG16 [41] 0.6607 0.66 /0.661 0.672 0.736
VGG16Inception [42] 0.7247 0.728/0.725 0.728 0.786
ViT [43] 0.7515 0.72/ 0.752 0.580 0.555
Swin-Transformer [44] 0.5110 0.51/ 0.511 0.52 0.619
Dinov2_Vitb14 [45] 0.667 0.665/0.667 0.678 0.742
ResNet50 [9] 0.6686 0.672 /0.669 0.677 0.74
DALAResNet50 0.8070 0.806/0.807 0.818 0.851
TABLE VIII: Parameters Count and Convergence Time of Network Models on BreakHis Dataset
@@
TABLE IX: Parameters Count and Convergence Time of Network Models on BACH Dataset
Network Models Total Parameters Average Time
SEResNet50 [39] 27074116 2921.67
DensNet121 [40] 9029996 3275.27
VGG16 [41] 134276934 3303.98
VGG16Inception [42] 38148134 3432.98
ViT [43] 51385352 2872.36
Swin-Transformer [44] 86753470 3488.72
Dinov2_Vitb14 [45] 86583556 2966.12
ResNet50 [9] 21805868 3723.49
DALAResNet50 24042692 2475.49
TABLE X: Parameters Count and Convergence Time of Network Models on Mini-DDSM Dataset
Network Models Total Parameters Average Time
SEResNet50 [39] 27073603 7272.75
DensNet121 [40] 9029483 6384.11
VGG16 [41] 134272835 7589.96
VGG16Inception [42] 38144035 8089.23
ViT [43] 51376131 6555.35
Swin-Transformer [44] 86746299 9715.44
Dinov2_Vitb14 [45] 86582787 6432.41
ResNet50 [9] 21797672 8268.68
DALAResNet50 24040643 5990.31
@@ IV-E Experiment metrics
Tables 5, 6, and 7 present the model’s performance on the binary classification BrekHis dataset, the four-class classification BACH dataset, and the three-class classification Mini-DDSM dataset. In these results, the model consistently outperforms nine other popular networks in terms of accuracy, F1 score, IBA, and GMean. This consistency highlights the significant advantages of the network and underscores its potential in medical image classification tasks.
The model was compared with eight other models in terms of parameter count and convergence speed on the binary classification BrekHe dataset, the four-class classification BACH dataset, and the three-class classification Mini-DDSM dataset, as shown in Tables 8, 9, and 10. Although DenseNet-121 has the fewest parameters, the proposed model achieved the shortest convergence time. This further reinforces the efficiency and robust potential of the proposed model in medical image classification, highlighting its superior performance and effectiveness.
@@ IV-F Ablation study
In this ablation study, we compared the performance of the DALAResNet50 model with the standard ResNet50 model on the binary classification BrekHis dataset at 40x, 100x, 200x, and 400x resolutions, as shown in Table 5 above. We also provided the confusion matrices for both models on the test set, as illustrated in Fig. 5. The results indicate that the DALAResNet50 model demonstrates higher sensitivity and specificity at all magnification levels, significantly reducing the number of false positives and false negatives. These findings suggest that the DALAResNet50 model outperforms the standard ResNet50 model in terms of classification accuracy and reliability.
@@ 1. Grad-CAM Calculation
2. Otsu Threshold Calculation
where:
  • $\sigma_{B}^{2}(T)$ is the between-class variance for threshold $T$.
@@ V-B Comparison between DT Grad-CAM and Grad-CAM Visualization
TABLE XI: Comparison between DT Grad-CAM and Grad-CAM
@@ VII DALAResNet50’s Future Therapeutic Impact and Integration into Existing Diagnostic Processes
Overall, the DALAResNet50 model, combined with DT Grad-CAM’s interpretability, holds great promise for transforming breast
@@
cancer diagnosis and treatment. Its integration into existing diagnostic workflows can lead to more precise, reliable, and efficient medical care, ultimately enhancing patient outcomes and advancing the field of histopathology.
@@ VIII Summary and conclusions
This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets. The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper’s model performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. Furthermore, this study introduces the DT Grad-CAM method to address the interpretability challenges often associated with deep learning models. DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, which provides clearer and more focused visualizations of significant regions. This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification. The integration of DT Grad-CAM with DALAResNet50 ensures high classification performance and enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis. In future research, we plan to apply the model presented in this paper to a broader range of medical fields.